OpenStudy (vanessalolololo):

Solve 4(x + 1) + 8 = 24.

OpenStudy (mathstudent55):

The first step is to get rid of the parentheses on the left side. Start by distributing the 4 on the left side.

OpenStudy (mathstudent55):

Can you do that and show it?

OpenStudy (zootedyoungin):

i hated that math when i was in 6th

OpenStudy (zootedyoungin):

2-4 step equations

OpenStudy (lapinblanc):

4(x+1)+8=24 multiply in 4 4x+4+8=24 combine like terms 4x+12=24 subtract 12 from both sides 4x=12 divide 4 from both sides to find x x=3

OpenStudy (mathstudent55):

The distributive property of multiplication over addition